OpenStudy (anonymous):

number patern,find the next two numbers. 1, 3, 9, 27, help thanks

OpenStudy (anonymous):

You mutiply the prior number by 3. So the next numbers are 81 and 243 because 27x3=81 and 81x3=243.

OpenStudy (anonymous):

that was very good thanks, my dad explained how you did that, but all my other questions were simple like 9, 18, 27, 36, and our class answer would be wrote like this +9 =18 +9=27 +9=27 to show the pattern any advice?

OpenStudy (anonymous):

All you have to do is show it as x3=27 x3=81 x 3 =243 (: If that's how you write it. Haha.